OpenStudy (anonymous):

i need step by step help of how to find the second derivative of -2/9x^7/9

OpenStudy (anonymous):

do you know how to calculate for the first derivative?

OpenStudy (anonymous):

the original problem was -x^2/9 so the derivative of that is the -2/9x^7/9

OpenStudy (anonymous):

\(\Large nx^{n-1}\\-x^{\frac{2}{9}}=> -\frac{2}{9}x^{\frac{2}{9}-1}=>-2/9x^{-7/9}\) then you should apply same rule: